OpenStudy (anonymous):

How Would I Find The Principal Square Root of 3025

OpenStudy (solomonzelman):

You know that \(\color{#000000 }{ \displaystyle 3025 }\) can be divided by 5 (because 3025 ends with a 5). \(\color{#000000 }{ \displaystyle 3025\div 5=? }\)

OpenStudy (solomonzelman):

((I am assuming you need an exact solution; not approximation))

OpenStudy (anonymous):

So The Principal Will BE 65

OpenStudy (solomonzelman):

not exactly, but you are one digit off

OpenStudy (solomonzelman):

(one digit is wrong)

OpenStudy (anonymous):

605

OpenStudy (solomonzelman):

Ok, please show your work...

OpenStudy (solomonzelman):

???

OpenStudy (solomonzelman):

that doesn't make sense, how do you just add a 5 outside there...

OpenStudy (anonymous):

i wasnt even done i thought i was dividing

OpenStudy (anonymous):

square root is a number multiplied by its self to get a perfect square

OpenStudy (mathmale):

I'd re-word that statement. The square root function is the INVERSE of the squaring function. The squaring function multiplies the input by itself. Example: 3^2 = 3*3 = 9.

OpenStudy (anonymous):

he said my own words

OpenStudy (mathmale):

Given 9, the Sqrt function takes you back to the original 3.

OpenStudy (mathmale):

You are given 3025 and are asked to find the principal square root. The result, when squared, will produce the square 3025. Is 65 a square root of 3025? To check, square 65.

OpenStudy (mathmale):

In the very old days, school kids were taught how to find square roots by hand (imagine!). Nowadays, calculators are commonly used for that. What approach does your teacher (or your learning materials) use?

OpenStudy (anonymous):

we use our phones

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 3025\div 5= 605 }\) \(\color{#000000 }{ \displaystyle 605\div 5= 121 }\) \(\color{#000000 }{ \displaystyle 121\div 11=11 }\) \(\color{#000000 }{ \displaystyle 11\div 11=1 }\) So, \(\color{#000000 }{ \displaystyle 3025=5\cdot 5 \cdot 11\cdot 11 }\)

OpenStudy (solomonzelman):

and yu had the correct solution in the begging by the way...

OpenStudy (solomonzelman):

So I just have to apologize; I apologize for what has happened.

OpenStudy (anonymous):

apologize to who